OpenStudy (anonymous):

Where am I going wrong?(Calc). THe question is: Find the volume of the triangle enclosed by the vertices (3,2)(3,4)(7,4) about the y-axis. What I have so far is the r(y)= the constant so 3. Then it tells me to make the equation of the line between the point (3,2)(7,4) with the formula y-y1/x-x1= y2-y1/x2-x1..which gave me 1/x-3=1. I plugged that into the integral and the answer is wrong. Can anyone help?

OpenStudy (jtvatsim):

what exactly does your integral look like that you tried? Example: Int from a to b of (whatever equation) dx

OpenStudy (anonymous):

top limit i have as 4, lower as 3

OpenStudy (anonymous):

then i was told from the lab to use the equation i get from that formula, but i have yet to get that right, then after that the constant squared, and the whole integral * pi

OpenStudy (anonymous):

i.e example (6y-4)^2-3^2 dy

OpenStudy (anonymous):

i can't seem to get the 6y-4(example) part correct

OpenStudy (jtvatsim):

ok, thank you for the details, that will help us give you a solution

OpenStudy (jtvatsim):

and just to be clear you are getting \[\frac{1}{x-3} = 1\] as your equation?

OpenStudy (anonymous):

yes. i have y as 3, so y-y1/x-x1 gave me 3-2/x-3, and y2-y1/x2-x1 gave me 4-2/4-2, so 1/x-3=1 is what im stuck with

OpenStudy (jtvatsim):

aha! there's the issue, you shouldn't be plugging in anything for y in the first part; you should not get 3 - 2

OpenStudy (jtvatsim):

that is an arbitrary y just like the x below it

OpenStudy (anonymous):

so 0-6/3-3= 0... and 1? 0 does not equal 1? lost

OpenStudy (jtvatsim):

If we collect the data this way it becomes more clear: (x1,y1) = (3,2) (x2,y2) = (7,4) The formula is (y - y1)/(x - x1) = (y2 - y1)/(x2 - x1) where y and x are unspecified to give you an actual equation

OpenStudy (jtvatsim):

You only need to plug in for x1, y1, x2, and y2.

OpenStudy (jtvatsim):

This should give you \[\frac{y - 2}{x - 3} = \frac{4 - 2}{7 - 3}\]

OpenStudy (jtvatsim):

Do you see that? :)

OpenStudy (anonymous):

yeah

OpenStudy (jtvatsim):

OK, so what would you think to do from here?

OpenStudy (anonymous):

y-2/x-3=1/2 then

OpenStudy (jtvatsim):

good, and then? :)

OpenStudy (anonymous):

if we want the line equation we need y=x so multiply the x-3 on both sides?

OpenStudy (jtvatsim):

ok, sounds good

OpenStudy (anonymous):

y= (1/2)(x-1)

OpenStudy (anonymous):

1/2x -1 is our equation?

OpenStudy (jtvatsim):

actually, hold up, is that supposed to be a minus inside the parentheses?

OpenStudy (anonymous):

where?

OpenStudy (jtvatsim):

We had y - 2 = (1/2)(x - 3), which becomes y - 2 = (1/2)x - 3/2, which becomes y = (1/2)x + 1/2 after adding 2 to both sides.

OpenStudy (jtvatsim):

it should be a plus sign is all I'm saying.

OpenStudy (anonymous):

so that gives my integral as: (pi) [integral] UPPER: 4 LOWER:3 -> (1/2y+1/2)^2-(3)^2 dy ?

OpenStudy (jtvatsim):

OK, got to be careful here, you can't just change the x's into y's like that. However, you are correct in trying to get only y's inside. Can you solve the equation of the line in terms of y's instead?

OpenStudy (jtvatsim):

How about taking y = (1/2)x + 1/2 and getting, y - 1/2 = (1/2)x, then, 2y - 1 = x. Does that make sense to you?

OpenStudy (anonymous):

oh crap. yeah

OpenStudy (anonymous):

so 82pi/3?

OpenStudy (jtvatsim):

lol, this is a tricky question... and the negative area should have scared you anyway. :)

OpenStudy (jtvatsim):

checking...

OpenStudy (anonymous):

hmm i did something wrong

OpenStudy (anonymous):

are my limits correct? 4/3?

OpenStudy (jtvatsim):

the integral should be \[\pi \int\limits_3^4 (2y-1)^2 - 3^2 dy\]

OpenStudy (anonymous):

thats what i have. isnt that 82/3? then with the pi 82pi/3?

OpenStudy (jtvatsim):

I'm almost there...

OpenStudy (jtvatsim):

yeah, that looks correct to me... is the answer in the book different?

OpenStudy (anonymous):

let me make sure i got all the numbers correct

OpenStudy (jtvatsim):

oh wait, the limits are 2 to 4 not 3!

OpenStudy (jtvatsim):

my bad, sorry on that

OpenStudy (jtvatsim):

because (3,2) has a y of 2 and (7,4) has a y of 4

OpenStudy (anonymous):

ha yeah thats it :)

OpenStudy (jtvatsim):

cool, awesome!

OpenStudy (anonymous):

got it right, thanks. 7th time is the wonder

OpenStudy (anonymous):

much appreciated.

OpenStudy (jtvatsim):

lol, no but seriously it takes like 10 times on these calc questions sometimes. :) Good job!

OpenStudy (anonymous):

cya in the future, thanks again

OpenStudy (jtvatsim):

take care